# Sequentially Auditing Differential Privacy

### v1

@@ C.1 Derivation of the alternative test
As mentioned in Section˜3.4, [51] studied a general class of e-processes of the form $W_{t}(\beta_{1}^{t})=\prod_{i\in[t]}(1+\beta_{i}(E_{i}-1))$, described in eq.˜4. Recall that for every $i\geq 1$, $\beta_{i}\in[0,1]$ and $E_{i}$ is an e-value for the null, meaning that $E_{i}$ is nonnegative and $\mathbb{E}_{H_{0}}[E_{i}]\leq 1$.
@@
Hence, the process $W_{t}^{\text{UP}}$ that results from sequentially choosing $\{\beta_{t}\}_{t\geq 0}$ with the Universal Portfolio (UP) algorithm [9] is a nonnegative supermartingale under the null. The regret guarantees of UP (Theorem˜E.3) give
@@ C.2 Statiscal properties of the alternative test
Part 2. We formally prove that there is a one-to-one map between the processes $\mathcal{K}_{t}(\lambda_{1}^{t})$ with $\lambda_{i}\in[0,1/(2+\tau)]$ for all $i\geq 1$ and $W_{t}(\beta_{1}^{t})$ with $\beta_{i}\in[0,1]$ for all $i\geq 1$. This follows from choosing $\beta_{i}=(2+\tau)\lambda_{i}$ for all $i\geq 1$:
$\mathcal{K}_{t}(\lambda_{1}^{t})$ $=\prod_{i\in[t]}\big{(}1+\lambda_{i}(f_{t}(X_{t})-f_{t}(Y_{t})-\tau)\big{)}$
@@
  • We are now able to optimize $\mathcal{K}_{t}(\lambda_{1}^{t})$ over $\lambda_{i}\in[0,1/(2+\tau)]$ for all $i\geq 1$. Recall that in order to use ONS we constrained to $\lambda_{i}\in[0,1/(4+2\tau)]$ so that the losses $\ell_{t}$ from Algorithm˜1 were Lipschitz.
@@
    $\max_{\lambda\in[0,1/(4+2\tau)]}\log(\mathcal{K}_{t}(\lambda))-\log(\mathcal{K}_{t}(\lambda_{1}^{t}))\leq 10\log(t)$
@@
    $\max_{\lambda\in[0,1/(2+\tau)]}\log(\mathcal{K}_{t}(\lambda))-\log(\mathcal{K}_{t}(\lambda_{1}^{t}))$ $=\max_{\beta\in[0,1]}\log(W_{t}(\beta))-\log(\tilde{W}_{t})$
@@ Appendix D Comparison to DP-Auditorium
In Section˜4.1, we reported the performance of sequential tests at detecting privacy violations of algorithms for private mean estimation (Table˜1) and commented on how these results compare to the ones reported in DP-Auditorium [26]. The presented sequential test differs from DP-Auditorium in many aspects, so in this section we decouple which of these aspects give the most improvement.
To make this comparison, we set $(\varepsilon,\delta)=(0.01,10^{-5})$ and report the rejection rates in Table˜3 over $10$ runs of different auditing algorithms that gradually move from our sequential algorithm to the $\mathrm{MMD}$-tester from [26]. Recall that the $\mathrm{MMD}$-tester constructs a confidence interval $[L,U]$ for $\mathrm{MMD}^{2}$ from a fixed batch of samples and then rejects if $L$ is greater than the square of the upper bound on the $\mathrm{MMD}$. Below, we list key differences between this algorithms and our sequential procedure.
  • Our algorithms use an improved bound on the $\mathrm{MMD}$ (see Theorem˜3.1), which is tighter than the one used in [26].
  • Both [26] and us use the RBF kernel. However, we use the median heuristic (MH) [17] to estimate the bandwidth using the first $20$ samples, which are then not used in the auditing procedure, while in DP-Auditorium the authors set the bandwidth to a constant. For Table˜3, we consider this constant to be $1$.
@@
In Table˜3, we evaluate the following auditing procedures. The algorithm ‘Batch’ is the $\mathrm{MMD}$-tester from [26]: it uses a fixed number of samples, the previous $\mathrm{MMD}$ bound from [26] and it fixes the bandwidth to $1$. The algorithm ‘Batch + new $\mathrm{MMD}$ bound’ replaces the $\mathrm{MMD}$ bound by the one that we obtained in Theorem˜3.1. The algorithm ‘Batch + MH’ modifies the $\mathrm{MMD}$-tester by implementing the median heuristic. Finally, ‘Batch + new $\mathrm{MMD}$ bound + MH’ implements the $\mathrm{MMD}$-tester along with the improved $\mathrm{MMD}$ bound and the median heuristic. Analogously, we define the sequential algorithms: ‘Sequential’ is an implementation of Algorithm˜1 that uses the $\mathrm{MMD}$ bound from [26] and does not use the median heurstic, and so on.
From Table˜3, we observe that some improvement comes from using a sequential algorithm [^3] Note that this improvement is not necessarily achieved by every sequential test. For example, one could construct a confidence sequence $([L_{t},U_{t}])$—a set of intervals such that with high probability, $L_{t}\leq\mathrm{MMD}^{2}\leq U_{t}$ for all $t\geq 1$— and reject if $L_{t}$ is larger than the $\mathrm{MMD}^{2}$ upper bound [32]. This procedure would typically not improve on the batch test. , while some of it also comes from using the median heuristic. We note that the improvement on the MMD bound is not very significant, since for the privacy parameters that we chose both bounds are actually close; our bound is better when $\varepsilon$ is large (e.g., more than $1$).
Table 3: We decouple the improvements of Algorithm˜1 over the $\mathrm{MMD}$-Tester from [26] by reporting detection rate of privacy violations with different auditing algorithms. We audit the privacy of three mechanisms defined in Section˜4: a $(0.01,10^{-5})$-DP mechanism (DPGaussian) and two non-DP mechanisms (NonDPGaussian1, NonDPGaussian2); we omit ‘Gaussian’ from their name in the table. We observe that some improvement comes from using our sequential algorithm, while some of it also comes from using the Median Heuristic (MH). We note that the improvement on the MMD bound is not very significant, since for the privacy parameters that we chose both bounds are actually close.
@@
Batch ($\mathrm{MMD}$-Tester [26]) 0 0 0.5
@@ E.2 Online Gradient Ascent
Theorem E.2 (Regret of OGA [36]).
@@ E.4 Universal Portfolio
starting from initial wealth $W_{0}(\beta)=1$. At time $t$, the universal portfolio strategy [9] defines a probability distribution over portfolios $\beta\in[0,1]$, weighted by their wealth up to time t1t-1. Specifically, under a prior density $\pi(\beta)$ on $[0,1]$, the time-$t$ portfolio is chosen as
@@
Theorem E.3 (Regret of UP (Theorem 3 in [10])).
@@ F.1 Perturbed uniform distributions
We begin by empirically demonstrating the behavior of our sequential test under two scenarios: when the null hypothesis of equal distributions is true and when it is false. For the former, we compare two 2-dimensional uniform distributions on the unit cube $0,1]^{2}$ where the auditing process $\mathcal{K}_{t}$ from Algorithm˜1 remains bounded below the rejection threshold $1/\alpha$ for a statistical confidence level $\alpha=0.05$. Figure 4 confirms that, as expected, we successfully control the Type I error at the desired statistical level when $H_{0}$ holds, across 100 simulations. For the latter scenario, we apply our test to compare a uniform distribution on $0,1]^{2}$ against a perturbed uniform distribution following the construction in [42] with one perturbation, across 100 simulations. In line with our theoretical predictions, the auditing process $\mathcal{K}_{t}$, which measures the cumulative evidence against the null hypothesis, grows exponentially under this alternative hypothesis, as illustrated in Figure 4. The test successfully rejects $H_{0}$ after, on average, collecting only 108 observations.
@@ F.2 Gaussian distributions
We then test our MMD-based sequential approach by comparing two Gaussian distributions while varying the dimensionality and separation between them. We fix one distribution as a $d$-dimensional standard normal distribution $\mathbb{P}_{1}\sim\mathcal{N}(0,I_{d})$, where $I_{d}$ is the $d$-dimensional identity matrix, varying $d$ from 1 to 5. Our test evaluates $H_{0}:\mathrm{MMD}(\mathbb{P}_{1},\mathbb{P}_{2})=0$ against $H_{a}:\mathrm{MMD}(\mathbb{P}_{1},\mathbb{P}_{2})>0$, with $\mathbb{P}_{2}\sim\mathcal{N}(\mu,I_{d})$ where we vary the center $\mu$ of the contrast distribution. We set the norm $\mu||$ to different values between 0 to $1$ to illustrate scenarios where the two distributions have different degrees of separation, making it progressively harder for the test to find evidence against the null as $\mu||$ is closer to 0. Note that when $\mu||=0$, the null hypothesis is true.
@@ Appendix G Auditing DP-SGD for Certain Privacy Regimes
Figure 7: Extended sequential audit results of non-private DP-SGD implementations over 2500 observations. The audit successfully rejects privacy hypotheses and establishes increasingly tight lower bounds on the privacy parameter.
@@ Appendix H Detailed Comparison with Related Work
Identifying worst-case datasets has been tackled through methods like grid search over datasets [11, 5], explicit construction under strong parametric assumptions on the mechanism [14], black-box optimization techniques that iteratively search for high-privacy-loss inputs [26], or the use of predefined canaries [33]. More recent approaches showed efficient testing relying on randomized canaries for specific mechanisms like the Gaussian mechanism in high dimensions [2]. Black-box optimization approaches treat divergence maximization as an objective but often lack formal guarantees on finding the true worst case [26]. Our proposed auditing method can be paired with any technique for identifying candidate worst-case dataset pairs.
@@
Estimating distances or divergences between distributions is a classical statistics problem [34, 47, 6]. Some methods employ optimization over function spaces, such as neural networks, to estimate $f$-divergences or Rényi divergences. While powerful, the finite-sample guarantees provided by some of these methods can depend on estimator properties (e.g., neural network architecture) and may become vacuous for DP auditing purposes [6]. Furthermore, many provide only asymptotic guarantees, whereas DP requires strict finite-sample bounds. Our work, in contrast, leverages sequential analysis to provide anytime-valid results with potentially much lower sample complexity.
Our approach relies on sequential testing with kernel methods, specifically the Maximum Mean Discrepancy (MMD)[19, 44], to construct test statistics. Prior work on privacy auditing has also used kernel methods, such as estimating regularized kernel Rényi divergence [13], but often requires strong assumptions (e.g., knowledge of covariance matrices) impractical in black-box settings or for mechanisms beyond Gaussian or Laplace.
Some auditing techniques require strong assumptions, like a discrete output space [18] or access to the mechanism’s internal randomness or probability density/mass functions [12]. Others need access to the cdf of the privacy loss random variable [14]. StatDP [11] requires semi-black-box access (e.g., running the mechanism without noise) [11]. While our method can incorporate white-box information to improve power, it fundamentally operates in the black-box setting. The applicability of these methods that require strong assumptions is limited in the blackbox setting.
For the general black-box setting, tools like DP-Sniper [5], DP-Opt [35], Delta-Siege [28], and Eureka [29] perform black-box testing by searching for an output event maximizing the probability difference between neighboring inputs. However, DP-Sniper is specific to pure $\varepsilon$-DP, Delta-Siege for $\rho$-ordered mechanisms" and all suffer from high sample complexity (millions of samples) in the batch setting.
A significant body of work focuses on auditing machine learning models, particularly those trained with DP-SGD. This includes membership inference attacks (MIA) [24, 37, 7, 23] and data reconstruction attacks [21, 4]. These methods often require white-box access to the model and sometimes large portions of the training data, but more importantly, work in the batch setting, requiring significant compute power. In the best case, they require training one full model end-to-end with canaries inserted every other step. For some easy to detect failures this might be wasting resources, and when the test fails, one has to start from scratch and previous samples cannot be reused for significant testing.

### v2

@@ C.1 Derivation of the alternative test
As mentioned in Section˜3.4, [waudby2025logopt] studied a general class of e-processes of the form $W_{t}(\beta_{1}^{t})=\prod_{i\in[t]}(1+\beta_{i}(E_{i}-1))$, described in eq.˜4. Recall that for every $i\geq 1$, $\beta_{i}\in[0,1]$ and $E_{i}$ is an e-value for the null, meaning that $E_{i}$ is nonnegative and $\mathbb{E}_{H_{0}}[E_{i}]\leq 1$.
@@
Hence, the process $W_{t}^{\text{UP}}$ that results from sequentially choosing $\{\beta_{t}\}_{t\geq 0}$ with the Universal Portfolio (UP) algorithm [cover1991up] is a nonnegative supermartingale under the null. The regret guarantees of UP (Theorem˜E.3) give
@@ C.2 Statiscal properties of the alternative test
Part 2. We formally prove that there is a one-to-one map between the processes $\mathcal{K}_{t}(\lambda_{t})$ with $\lambda_{i}\in[0,1/(2+\tau)]$ for all $i\geq 1$ and $W_{t}(\beta_{1}^{t})$ with $\beta_{i}\in[0,1]$ for all $i\geq 1$. This follows from choosing $\beta_{i}=(2+\tau)\lambda_{i}$ for all $i\geq 1$:
$\mathcal{K}_{t}$ $=\prod_{i\in[t]}\big(1+\lambda_{i}(f_{t}(X_{t})-f_{t}(Y_{t})-\tau)\big)$
@@
  • We are now able to optimize $\mathcal{K}_{t}$ over $\lambda_{i}\in[0,1/(2+\tau)]$ for all $i\geq 1$. Recall that in order to use ONS we constrained to $\lambda_{i}\in[0,1/(4+2\tau)]$ so that the losses $\ell_{t}$ from Algorithm˜1 were Lipschitz.
@@
    $\max_{\lambda\in[0,1/(4+2\tau)]}\log(\mathcal{K}_{t}(\lambda))-\log(\mathcal{K}_{t})\leq 10\log(t)$
@@
    $\max_{\lambda\in[0,1/(2+\tau)]}\log(\mathcal{K}_{t}(\lambda))-\log(\mathcal{K}_{t})$ $=\max_{\beta\in[0,1]}\log(W_{t}(\beta))-\log(\tilde{W}_{t})$
@@ Appendix D Comparison to DP-Auditorium
In Section˜4.1, we reported the performance of sequential tests at detecting privacy violations of algorithms for private mean estimation (Table˜1) and commented on how these results compare to the ones reported in DP-Auditorium [2024DPAuditorium]. The presented sequential test differs from DP-Auditorium in many aspects, so in this section we decouple which of these aspects give the most improvement.
To make this comparison, we set $(\varepsilon,\delta)=(0.01,10^{-5})$ and report the rejection rates in Table˜3 over $10$ runs of different auditing algorithms that gradually move from our sequential algorithm to the $\mathrm{MMD}$-tester from [2024DPAuditorium]. Recall that the $\mathrm{MMD}$-tester constructs a confidence interval $[L,U]$ for $\mathrm{MMD}^{2}$ from a fixed batch of samples and then rejects if $L$ is greater than the square of the upper bound on the $\mathrm{MMD}$. Below, we list key differences between this algorithms and our sequential procedure.
  • Our algorithms use an improved bound on the $\mathrm{MMD}$ (see Theorem˜3.1), which is tighter than the one used in [2024DPAuditorium].
  • Both [2024DPAuditorium] and us use the RBF kernel. However, we use the median heuristic (MH) [garreau2017large] to estimate the bandwidth using the first $20$ samples, which are then not used in the auditing procedure, while in DP-Auditorium the authors set the bandwidth to a constant. For Table˜3, we consider this constant to be $1$.
@@
In Table˜3, we evaluate the following auditing procedures. The algorithm ‘Batch’ is the $\mathrm{MMD}$-tester from [2024DPAuditorium]: it uses a fixed number of samples, the previous $\mathrm{MMD}$ bound from [2024DPAuditorium] and it fixes the bandwidth to $1$. The algorithm ‘Batch + new $\mathrm{MMD}$ bound’ replaces the $\mathrm{MMD}$ bound by the one that we obtained in Theorem˜3.1. The algorithm ‘Batch + MH’ modifies the $\mathrm{MMD}$-tester by implementing the median heuristic. Finally, ‘Batch + new $\mathrm{MMD}$ bound + MH’ implements the $\mathrm{MMD}$-tester along with the improved $\mathrm{MMD}$ bound and the median heuristic. Analogously, we define the sequential algorithms: ‘Sequential’ is an implementation of Algorithm˜1 that uses the $\mathrm{MMD}$ bound from [2024DPAuditorium] and does not use the median heurstic, and so on.
From Table˜3, we observe that some improvement comes from using a sequential algorithm [^3] Note that this improvement is not necessarily achieved by every sequential test. For example, one could construct a confidence sequence $([L_{t},U_{t}])$—a set of intervals such that with high probability, $L_{t}\leq\mathrm{MMD}^{2}\leq U_{t}$ for all $t\geq 1$— and reject if $L_{t}$ is larger than the $\mathrm{MMD}^{2}$ upper bound [manole2023martingale]. This procedure would typically not improve on the batch test. , while some of it also comes from using the median heuristic. We note that the improvement on the MMD bound is not very significant, since for the privacy parameters that we chose both bounds are actually close; our bound is better when $\varepsilon$ is large (e.g., more than $1$).
Table 3: We decouple the improvements of Algorithm˜1 over the $\mathrm{MMD}$-Tester from [2024DPAuditorium] by reporting detection rate of privacy violations with different auditing algorithms. We audit the privacy of three mechanisms defined in Section˜4: a $(0.01,10^{-5})$-DP mechanism (DPGaussian) and two non-DP mechanisms (NonDPGaussian1, NonDPGaussian2); we omit ‘Gaussian’ from their name in the table. We observe that some improvement comes from using our sequential algorithm, while some of it also comes from using the Median Heuristic (MH). We note that the improvement on the MMD bound is not very significant, since for the privacy parameters that we chose both bounds are actually close.
@@
Batch ($\mathrm{MMD}$-Tester [2024DPAuditorium]) 0 0 0.5
@@ E.2 Online Gradient Ascent
Theorem E.2 (Regret of OGA [orabona2023onlinelearning]).
@@ E.4 Universal Portfolio
starting from initial wealth $W_{0}(\beta)=1$. At time $t$, the universal portfolio strategy [cover1991up] defines a probability distribution over portfolios $\beta\in[0,1]$, weighted by their wealth up to time t1t-1. Specifically, under a prior density $\pi(\beta)$ on $[0,1]$, the time-$t$ portfolio is chosen as
@@
Theorem E.3 (Regret of UP (Theorem 3 in [cover2002up_sideinfo])).
@@ F.1 Perturbed uniform distributions
We begin by empirically demonstrating the behavior of our sequential test under two scenarios: when the null hypothesis of equal distributions is true and when it is false. For the former, we compare two 2-dimensional uniform distributions on the unit cube $0,1]^{2}$ where the auditing process $\mathcal{K}_{t}$ from Algorithm˜1 remains bounded below the rejection threshold $1/\alpha$ for a statistical confidence level $\alpha=0.05$. Figure 4 confirms that, as expected, we successfully control the Type I error at the desired statistical level when $H_{0}$ holds, across 100 simulations. For the latter scenario, we apply our test to compare a uniform distribution on $0,1]^{2}$ against a perturbed uniform distribution following the construction in [schrab2023mmd] with one perturbation, across 100 simulations. In line with our theoretical predictions, the auditing process $\mathcal{K}_{t}$, which measures the cumulative evidence against the null hypothesis, grows exponentially under this alternative hypothesis, as illustrated in Figure 4. The test successfully rejects $H_{0}$ after, on average, collecting only 108 observations.
@@ F.2 Gaussian distributions
We then test our MMD-based sequential approach by comparing two Gaussian distributions while varying the dimensionality and separation between them. We fix one distribution as a $d$-dimensional standard normal distribution $\mathbb{P}_{1}\sim\mathcal{N}(0,I_{d})$, where $I_{d}$ is the $d$-dimensional identity matrix, varying $d$ from 1 to 5. Our test evaluates $H_{0}:\mathrm{MMD}(\mathbb{P}_{1},\mathbb{P}_{2})=0$ against $H_{1}:\mathrm{MMD}(\mathbb{P}_{1},\mathbb{P}_{2})>0$, with $\mathbb{P}_{2}\sim\mathcal{N}(\mu,I_{d})$ where we vary the center $\mu$ of the contrast distribution. We set the norm $\mu||$ to different values between 0 to $1$ to illustrate scenarios where the two distributions have different degrees of separation, making it progressively harder for the test to find evidence against the null as $\mu||$ is closer to 0. Note that when $\mu||=0$, the null hypothesis is true.
@@ Appendix G Auditing DP-SGD for Certain Privacy Regimes
Figure 7: Extended sequential audit results of non-private DP-SGD implementations over 2,500 observations. The audit successfully rejects privacy hypotheses and establishes increasingly tight lower bounds on the privacy parameter.
@@ Appendix H Detailed Comparison with Related Work
Identifying worst-case datasets has been tackled through methods like grid search over datasets [StatDP, dpsniper], explicit construction under strong parametric assumptions on the mechanism [DGKK22], black-box optimization techniques that iteratively search for high-privacy-loss inputs [2024DPAuditorium], or the use of predefined canaries [2023tightauditingDPML]. More recent approaches showed efficient testing relying on randomized canaries for specific mechanisms like the Gaussian mechanism in high dimensions [AKOOMS23]. Black-box optimization approaches treat divergence maximization as an objective but often lack formal guarantees on finding the true worst case [2024DPAuditorium]. Our proposed auditing method can be paired with any technique for identifying candidate worst-case dataset pairs.
@@
Estimating distances or divergences between distributions is a classical statistics problem [NWJ10, sriperumbudur2012empirical, BDKRW21]. Some methods employ optimization over function spaces, such as neural networks, to estimate $f$-divergences or Rényi divergences. While powerful, the finite-sample guarantees provided by some of these methods can depend on estimator properties (e.g., neural network architecture) and may become vacuous for DP auditing purposes [BDKRW21]. Furthermore, many provide only asymptotic guarantees, whereas DP requires strict finite-sample bounds. Our work, in contrast, leverages sequential analysis to provide anytime-valid results with potentially much lower sample complexity.
Our approach relies on sequential testing with kernel methods, specifically the Maximum Mean Discrepancy (MMD)[gretton2012kernel, shekhar2023nonparametric2sampletesting], to construct test statistics. Prior work on privacy auditing has also used kernel methods, such as estimating regularized kernel Rényi divergence [DM22], but often requires strong assumptions (e.g., knowledge of covariance matrices) impractical in black-box settings or for mechanisms beyond Gaussian or Laplace.
Some auditing techniques require strong assumptions, like a discrete output space [GM18] or access to the mechanism’s internal randomness or probability density/mass functions [DJT13]. Others need access to the cdf of the privacy loss random variable [DGKK22]. StatDP [StatDP] requires semi-black-box access (e.g., running the mechanism without noise) [StatDP]. While our method can incorporate white-box information to improve power, it fundamentally operates in the black-box setting. The applicability of these methods that require strong assumptions is limited in the blackbox setting.
For the general black-box setting, tools like DP-Sniper [dpsniper], DP-Opt [niu2022dp], Delta-Siege [lokna2023group], and Eureka [lu2022eureka] perform black-box testing by searching for an output event maximizing the probability difference between neighboring inputs. However, DP-Sniper is specific to pure $\varepsilon$-DP, Delta-Siege for $\rho$-ordered mechanisms" and all suffer from high sample complexity (millions of samples) in the batch setting.
A significant body of work focuses on auditing machine learning models, particularly those trained with DP-SGD. This includes membership inference attacks (MIA) [JE19, ROF21, CYZF20, jagielski2020auditing] and data reconstruction attacks [Guo22, Balle22]. These methods often require white-box access to the model and sometimes large portions of the training data, but more importantly, work in the batch setting, requiring significant compute power. In the best case, they require training one full model end-to-end with canaries inserted every other step. For some easy to detect failures this might be wasting resources, and when the test fails, one has to start from scratch and previous samples cannot be reused for significant testing.